# Prioritized Generative Replay

### v1

@@ Abstract
Sample-efficient online reinforcement learning often uses replay buffers to store experience for reuse when updating the value function. However, uniform replay is inefficient, since certain classes of transitions can be more relevant to learning. While prioritization of more useful samples is helpful, this strategy can also lead to overfitting, as useful samples are likely to be more rare. In this work, we instead propose a prioritized, parametric version of an agent’s memory, using generative models to capture online experience. This paradigm enables (1) densification of past experience, with new generations that benefit from the generative model’s generalization capacity and (2) guidance via a family of “relevance functions” that push these generations towards more useful parts of an agent’s acquired history. We show this recipe can be instantiated using conditional diffusion models and simple relevance functions such as curiosity- or value-based metrics. Our approach consistently improves performance and sample efficiency in both state- and pixel-based domains. We expose the mechanisms underlying these gains, showing how guidance promotes diversity in our generated transitions and reduces overfitting. We also showcase how our approach can train policies with even higher update-to-data ratios than before, opening up avenues to better scale online RL agents.
@@ 1 Introduction
One central problem of online reinforcement learning (RL) lies in extracting signal from a continuous stream of experience. Not only does the non-i.i.d. form of this data induce learning instabilities, but agents may lose out on near-term experience that is not immediately useful, but becomes important much later. The standard solution to these problems is to use a replay buffer Lin (1992); Mnih et al. (2015) as a form of memory. By storing a dataset of transitions to enable batch-wise sampling, the algorithm can decorrelate online observations and revisit past experience later in training.
@@
An ideal relevance function $\mathcal{F}$ should be easy to compute, and naturally identify the most relevant experience. Intuitively, we should generate transitions at the “frontier” of the agent’s experience. These are regions of transition space that our generative model can accurately capture from the agent’s memory, but our policy has not yet completely mastered. We investigate a series of possible functions, concluding that intrinsic curiosity reliably approximates this distribution.
@@ 2 Related Work
RL from synthetic data. The use of generative training data has a long history in reinforcement learning (Shin et al., 2017; Raghavan et al., 2019; Imre, 2021; Hafner et al., 2019a). But only with the recent advent of powerful diffusion models have such methods achieved parity with methods trained on real data (Janner et al., 2022; Ajay et al., 2022; Zhu et al., 2023). The use of diffusion for RL was first introduced by Diffuser (Janner et al., 2022). The authors propose an offline diffusion model for generating trajectories of states and actions, and directly generate plans which reach goal states or achieve high rewards. Decision Diffuser (Ajay et al., 2022) generates state-only trajectories, and employs an inverse kinematics model to generate corresponding actions. More recently, works like (Ding et al., 2024; He et al., 2024) leverage diffusion models to augment datasets for offline RL, improving training stability. Most similar to our approach is synther (Lu et al., 2024), which augments the replay buffer online and can be seen as an unguided form of our framework. In contrast to these previous works, we posit that the strength of synthetic data models is that they can be guided towards novel transitions that are off-policy yet more relevant for learning.
@@ 3 Background
$\mathbb{E}_{x^{0}\sim\mathcal{D},\epsilon\sim\mathcal{N}(0,\mathbf{I}),y,n\sim\textrm{Unif}(1,N),p\sim\textrm{Bernoulli}\left(p_{\textrm{uncond}}\right)}\lVert\epsilon_{\theta}\left(x^{n},n,(1-p)\cdot y+p\cdot\emptyset\right)\rVert^{2}_{2},$ (2)
@@ 4 Prioritized Generative Replay
In this section, we introduce Prioritized Generative Replay (pgr). At its core, pgr involves a parametric generative replay buffer that can be guided by a variety of relevance criteria. We first provide intuition and motivation for such a framework, and concretize how it can be instantiated. Next, we compare and contrast between various instantiations of relevance functions. We argue that we should use functions which promote generating higher-novelty states as they implicitly increase the diversity of relevant transitions our replay buffer, leading to more sample-efficient learning.
@@ 4.1 Motivation
  • 1. $|\mathcal{D}|$ must be sufficiently large and diverse to prevent overfitting (hard to guarantee online)
@@ 4.2 Relevance Functions
Return-as-relevance sensibly pushes generations to be more on-policy, since $\pi$ by construction seeks out high $Q$-estimates states. Also, many online RL algorithms already learn a $Q$-function, and so we readily satisfy the first condition (Mnih et al., 2015). However, the second condition is not adequately addressed by this choice of $\mathcal{F}$. In particular, the diversity of high-return transitions might in practice be quite low, making overfitting to the conditional generations more likely.
@@
We argue that a natural way to reduce overfitting is via some relevance function which promotes generation diversity. As prior work has shown, an effective way to decrease overfitting to early, noisy signal in online RL is to leverage diverse experience (Zhang et al., 2018). To achieve this diversity, we model $\mathcal{F}$ after exploration objectives which promote engaging with “higher-novelty” transitions that are more rarely seen (Strehl & Littman, 2008). Moreover, by learning a separate function entirely, we decorrelate our relevance function from the $Q$-function, making overfitting less likely.
@@ 5 Experiments
Model details, training and evaluation. For state-based tasks, in additional to the sac and redq model-free baselines, we also include two strong model-based online RL baselines in mbpo (Janner et al., 2019) and dreamer-v3 (Hafner et al., 2023). We also compare against synther (Lu et al., 2024), a recent work which learns an unconditional generative replay model, allowing sac to be trained with an update-to-data (UTD) ratio of 20. To facilitate conditional sampling, during training we randomly discard the scalar given by our relevance function $\mathcal{F}$ with probability 0.25.
@@ 5.1 Results
As shown in Table 1 and Table 3, all variants of pgr successfully solve the tasks, with curiosity guidance consistently outperforming both model-free (sac, redq and drq-v2), and model-based (mbpo and dreamer-v3) algorithms, as well as the unconditional generation baseline (synther).
Comparison to prioritized experience replay (PER). We also compare pgr to PER baselines that use different measures of priority. In particular, we train redq with a prioritized replay buffer, using the classic TD-error (Schaul et al., 2015), or a priority function based on Eq. 5. The former is comparable to our pgr approach using Eq. 4 as relevance, and the latter to using Eq. 5. As shown in  Fig. 3a, pgr demonstrates superior performance in both cases. This emphasizes the importance of densifying the replay distribution with generations, and not simply reweighting past experience.
@@
Comparison to exploration bonuses. We examine how baselines improve when given a bonus in the form of an intrinsic curiosity reward (c.f. Eq. 5). In Fig. 3b, we see that curiosity-pgr continues to outperform synther or redq when either is augmented this way. This suggests pgr goes beyond just improving exploration. We propose our gains are a result of generating higher novelty transitions from the replay buffer with higher diversity. This additional diversity thereby reduces overfitting of the $Q$-function to synthetic transitions. We provide further evidence for this idea in Section 5.2.
@@ 5.2 Sample Efficiency
In this section, we show pgr achieves its strong performance in a sample-efficient manner, and reveal the mechanisms underlying this sample-efficiency. As seen in Fig. 4a, for tasks with state-based inputs, synther often attains its best performance around $100K$ environment steps. In contrast, curiosity-pgr is able to match this performance in both the cheetah-run and quadruped-walk tasks after only ${\sim}50K$ steps. Especially noteworthy is the finger-turn-hard task, where synther actually underperforms compared to the vanilla model-free redq baseline, while our curiosity-pgr continues outperforming. We argue conditioning is particularly valuable in sparse reward tasks, where efficient exploration of the environment presents a significant challenge. These observations hold for pixel-based tasks. As we see in Fig. 4b, while synther is eventually overtaken by drq-v2 in both environments, our curiosity-pgr continues to consistently improve over drq-v2.
@@
Moreover, naïve choices for conditioning fail entirely. For example, training on a larger quantity of higher reward transitions should intuitively yield a high-reward policy. However, as we show in  Fig. 4 and Table 1, naïvely conditioning on high reward (Reward-pgr) actually results in worse performance than any of the other pgr variants we propose, and in fact does worse than unconditional generation (synther). This is because greedily densifying high reward transitions does not actually provide the policy with any data on how to navigate to those high reward environment regions. Thus, it is non-trivial to even know what the correct conditioning signal should be to obtain strong performance.
Figure 6: Curiosity-pgr reduces overfitting and improves diversity of replay data. (a) Dormant ratio (Xu et al., 2023) (DR) of policy networks for different approaches. DR is consistently lower for pgr, indicating a minimally overfitting policy. (b) Curiosity $\mathcal{F}$-values throughout training for unconditional baseline and pgr. We show that the distribution of states accessed by the curiosity-pgr policy is significantly shifted towards higher novelty environment transitions.
@@ 6 Conclusion
In this work, we propose Prioritized Generative Replay (pgr): a parametric, prioritized formulation of an online agent’s memory based on an end-to-end learned conditional generative model. pgr conditions on a relevance function $\mathcal{F}$ to guide generation towards more learning-informative transitions, improving sample efficiency in both state- and pixel-based tasks. We show that it is not conditional generation itself, but rather conditioning on the correct $\mathcal{F}$ that is critical. We identify curiosity as the right choice for $\mathcal{F}$, and provide a decisive answer for why: curiosity-pgr improves the diversity of generative replay, thereby reducing overfitting to synthetic data. pgr also showcases a promising formula for scalable training using synthetic data, opening up new directions in generative RL.

### v2

@@ Abstract
Sample-efficient online reinforcement learning often uses replay buffers to store experience for reuse when updating the value function. However, uniform replay is inefficient, since certain classes of transitions can be more relevant to learning. While prioritization of more useful samples is helpful, this strategy can also lead to overfitting, as useful samples are likely to be more rare. In this work, we instead propose a prioritized, parametric version of an agent’s memory, using generative models to capture online experience. This paradigm enables (1) densification of past experience, with new generations that benefit from the generative model’s generalization capacity and (2) guidance via a family of “relevance functions” that push these generations towards more useful parts of an agent’s acquired history. We show this recipe can be instantiated using conditional diffusion models and simple relevance functions such as curiosity- or value-based metrics. Our approach consistently improves performance and sample efficiency in both state- and pixel-based domains. We expose the mechanisms underlying these gains, showing how guidance promotes diversity in our generated transitions and reduces overfitting. We also showcase how our approach can train policies with even higher update-to-data ratios than before, opening up avenues to better scale online RL agents. [^1] Project page available at: https://pgenreplay.github.io
@@ 1 Introduction
A central problem in online reinforcement learning (RL) involves extracting signal from a continuous stream of experience. Not only does the non-i.i.d. form of this data induce learning instabilities, but agents may lose out on near-term experience that is not immediately useful, but becomes important much later. The standard solution to these problems is to use a replay buffer as a form of memory (Lin, 1992; Mnih et al., 2015). By storing a dataset of transitions to enable batch-wise sampling, the algorithm can decorrelate online observations and revisit past experience later in training.
@@
An ideal relevance function $\mathcal{F}$ should be easy to compute, and naturally identify the most relevant experience. Intuitively, we should generate transitions at the “frontier” of the agent’s experience. These are regions of transition space that our generative model can accurately capture from the agent’s memory, but our policy has not yet completely mastered. We investigate a series of possible functions, concluding that intrinsic curiosity can successfully approximate this distribution.
@@ 2 Related Work
RL from synthetic data. The use of generative training data has a long history in reinforcement learning (Shin et al., 2017; Raghavan et al., 2019; Imre, 2021; Hafner et al., 2019a). But only with the recent advent of powerful diffusion models have such methods achieved parity with methods trained on real data (Janner et al., 2022; Ajay et al., 2022; Zhu et al., 2023). The use of diffusion for RL was first introduced by Diffuser (Janner et al., 2022). The authors propose an offline diffusion model for generating trajectories of states and actions, and directly generate plans which reach goal states or achieve high rewards. Decision Diffuser (Ajay et al., 2022) generates state-only trajectories, and employs an inverse kinematics model to generate corresponding actions. More recently, works like Ding et al. (2024); He et al. (2024) leverage diffusion models to augment datasets for offline RL, improving training stability. Most similar to our approach is synther (Lu et al., 2024), which augments the replay buffer online and can be seen as an unguided form of our framework. In contrast to these previous works, we posit that the strength of synthetic data models is that they can be guided towards novel transitions that are off-policy yet more relevant for learning.
@@ 3 Background
$\mathbb{E}_{x^{0}\sim\mathcal{D},\epsilon\sim\mathcal{N}(0,\mathbf{I}),y,n\sim\textrm{Unif}(1,N),p\sim\textrm{Bernoulli}\left(p_{\textrm{uncond}}\right)}\left\lVert\epsilon_{\theta}\left(x^{n},n,(1-p)\cdot y+p\cdot\emptyset\right)\right\rVert^{2}_{2},$ (2)
@@ 4 Prioritized Generative Replay
In this section, we introduce Prioritized Generative Replay (pgr). At its core, pgr involves a parametric generative replay buffer that can be guided by a variety of relevance criteria. We first provide intuition and motivation for such a framework, and concretize how it can be instantiated. Next, we compare and contrast between various instantiations of relevance functions. Most interestingly, we will empirically show that a good default choice for sample-efficient learning is intrinsic curiosity.
@@ 4.1 Motivation
  • 1. $\mathcal{D}$ must be sufficiently large and diverse to prevent overfitting (hard to guarantee online)
@@ 4.2 Relevance Functions
Return-as-relevance sensibly pushes generations to be more on-policy, since $\pi$ by construction seeks out high $Q$-estimate states. Also, many online RL algorithms already learn a $Q$-function, and so we readily satisfy the first condition (Mnih et al., 2015). However, the second condition is not adequately addressed by this choice of $\mathcal{F}$. In particular, the diversity of high-return transitions might in practice be quite low, making overfitting to the conditional generations more likely.
@@
Naturally then, we might consider reducing overfitting via some relevance function which promotes generation diversity. As prior work has shown, an effective way to decrease overfitting to early, noisy signal in online RL is to leverage diverse experience (Zhang et al., 2018). To achieve this diversity, we model $\mathcal{F}$ after exploration objectives which promote engaging with “higher-novelty” transitions that are more rarely seen (Strehl & Littman, 2008). Moreover, by learning a separate function entirely, we decorrelate our relevance function from the $Q$-function, making overfitting less likely.
@@ 5 Experiments
Model details, training and evaluation. For state-based tasks, in addition to the sac and redq model-free baselines, we also include two strong model-based online RL baselines in mbpo (Janner et al., 2019) and dreamer-v3 (Hafner et al., 2023). We also compare against synther (Lu et al., 2024), a recent work which learns an unconditional generative replay model, allowing sac to be trained with an update-to-data (UTD) ratio of 20. To facilitate conditional sampling, during training we randomly discard the scalar given by our relevance function $\mathcal{F}$ with probability 0.25.
@@ 5.1 Results
As shown in Table 1 and Table 3, all variants of pgr successfully solve the tasks, with curiosity guidance consistently outperforming both model-free (sac, redq and drq-v2), and model-based (mbpo and dreamer-v3) algorithms, as well as unconditional generation (synther). We also investigate alternative measures of curiosity such as random network distillation (Burda et al., 2018) and episodic curiosity (Savinov et al., 2018) in stochastic environments, with results in Appendix A.
Comparison to prioritized experience replay (PER). We also compare pgr to PER baselines that use different measures of priority. In particular, we train redq with a prioritized replay buffer, using the classic TD-error (Schaul et al., 2015), or a priority function based on Eq. 5. The former is compared to our pgr approach using Eq. 4 as relevance, and the latter to using Eq. 5. As shown in  Fig. 3a, pgr demonstrates superior performance in both cases. This emphasizes the importance of densifying the replay distribution with generations, and not simply reweighting past experience.
@@
Comparison to exploration bonuses. We examine how baselines improve when given a bonus in the form of an intrinsic curiosity reward (c.f. Eq. 5). In Fig. 3b, we see that curiosity-pgr continues to outperform synther or redq when either is augmented this way. This suggests pgr goes beyond just improving exploration. We provide additional evidence for this claim in Appendix B. We propose our gains are a result of generating higher novelty transitions from the replay buffer, with higher diversity. This additional diversity thereby reduces overfitting of the $Q$-function to synthetic transitions. We provide further evidence for this idea in Section 5.2.
@@ 5.2 Sample Efficiency
[…]
@@
Moreover, naïve choices for conditioning fail entirely. For example, training on a larger quantity of higher reward transitions should intuitively yield a high-reward policy. However, as we show in  Fig. 4 and Table 1, naïvely conditioning on high reward (Reward-pgr) actually results in worse performance than any of the other pgr variants we propose, and in fact does worse than unconditional generation (synther). This is because greedily densifying high reward transitions does not actually provide the policy with any data on how to navigate to those high reward environment regions. Thus, for obtaining strong performance, it is important to be thoughtful in our choice of $\mathcal{F}$.
Figure 6: Curiosity-pgr reduces overfitting and improves diversity of replay data. (a) Dormant ratio (Xu et al., 2023) (DR) of policy networks for different approaches. DR is consistently lower for pgr, indicating a minimally overfitting policy. (b) Curiosity $\mathcal{F}$-values throughout training for the unconditional baseline and pgr. We show that the distribution of states accessed by the curiosity-pgr policy is significantly shifted towards higher novelty environment transitions.
@@ 6 Conclusion
In this work, we propose Prioritized Generative Replay (pgr): a parametric, prioritized formulation of an online agent’s memory based on an end-to-end learned conditional generative model. pgr conditions on a relevance function $\mathcal{F}$ to guide generation towards more learning-informative transitions, improving sample efficiency in both state- and pixel-based tasks. We show that it is not conditional generation itself, but rather conditioning on the correct $\mathcal{F}$ that is critical. We identify curiosity as a good default choice for $\mathcal{F}$, and decisively answer why: curiosity-pgr improves the diversity of generative replay, and reduces overfitting to synthetic data. pgr also showcases a promising new formula for scalable training with synthetic data, opening up new directions in generative RL.